# Empirical Bayes for Data integration

### v1

@@ 5.2. Expectation-Maximization algorithm
Consider an EM algorithm setting where the likelihood for the full data $(\bm{y},\bm{\gamma})$ is given by $p(\bm{y},\bm{\gamma}\mid\bm{\omega})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma}\mid\bm{\omega})$, and where $\bm{\gamma}$ are the latent variables in the usual definition of the EM algorithm. The algorithm requires an E-step where one obtains the expected full data log-likelihood with respect to $\bm{\gamma}$, given the data $\bm{y}$ and the current parameter estimate $\hat{\bm{\omega}}^{(k)}$. This requires obtaining $\pi(\bm{\gamma}\mid\bm{y},\hat{\bm{\omega}}^{(k)})$ and, if an exact computation is too costly (one cannot enumerate all models), one may use an stochastic EM algorithm where $\pi(\bm{\gamma}\mid\bm{y},\hat{\bm{\omega}}^{(k)})$ is replaced by the frequency of posterior samples visiting model $\bm{\gamma}$. Given such $\pi(\bm{\gamma}\mid\bm{y},\hat{\bm{\omega}}^{(k)})$, the M-step seeks to maximize
@@
$\nabla_{\bm{\omega}}f(\bm{\omega})=\sum_{j=1}^{p}\frac{\nabla_{\bm{\omega}}m_{j}(\bm{\omega})}{m_{j}(\bm{\omega})[1-m_{j}(\bm{\omega})]}\left[\pi(\gamma_{j}=1\mid\bm{y},\hat{\bm{\omega}})-\pi(\gamma_{j}=1\mid\bm{\omega})\right].$
In the particular case $m_{j}(\bm{\omega})=1/(1+e^{\bm{z}_{j}^{T}\bm{\omega}})$, then $\nabla_{\bm{\omega}}f(\bm{\omega})=\bm{Z}^{T}[\hat{\bm{\pi}}-{\bf m}(\bm{\omega})]$ and $\nabla_{\bm{\omega}}^{2}f(\bm{\omega})=-\bm{Z}^{T}\bm{D}(\bm{\omega})\bm{Z}$ where $\hat{\bm{\pi}}=(\pi(\gamma_{1}\mid\bm{y},\hat{\bm{\omega}}),\ldots,\pi(\gamma_{p}\mid\bm{y},\hat{\bm{\omega}}))^{T}$, ${\bf m}(\bm{\omega})=(m_{1}(\bm{\omega}),\ldots,m_{p}(\bm{\omega}))^{T}$ and $\bm{D}(\bm{\omega})=\mbox{diag}(m_{1}(\bm{\omega}),\ldots,m_{p}(\bm{\omega}))$.
[…]
@@
Let $\bm{U}$ be the sub-matrix containing the unique rows of the $p\times q$ matrix $\bm{Z}$. Suppose that $\bm{U}$ has $q$ rows and rank $q$, $m_{j}(\bm{\omega})=1/(1+e^{-\bm{z}_{j}^{T}\bm{\omega}})$, and $\bm{\omega}\sim N({\bf 0},g(\bm{Z}^{T}\bm{Z}/p)^{-1})$. Then $\nabla_{\bm{\omega}}f(\hat{\bm{\omega}})+\nabla_{\bm{\omega}}\log\pi(\hat{\bm{\omega}})={\bf 0}$, where $\hat{\bm{\omega}}=\bm{U}^{-1}\tilde{\bm{\omega}}$ and $\tilde{\omega}_{j}=h(a_{j},gp)$, where $h(a,c)$ is the solution to $1/(1+e^{-w})+w/c=a$, $a_{j}=\sum_{i=1}^{p}\tilde{z}_{ij}^{T}\hat{\pi}_{i}/\sum_{i=1}^{p}z_{ij}$, $\tilde{\bm{Z}}=\bm{Z}\bm{U}^{-1}$ and $\hat{\bm{\pi}}$ is as in Corollary 1.
@@ 5.3. Other computational methods
Upon convergence, $\bm{\omega}^{(l)}$ is by definition a sample from $\pi(\bm{\omega}\mid\bm{y})$. One may set $\hat{\bm{\omega}}$ to the posterior mean $L^{-1}\sum_{l=1}^{L}\bm{\omega}^{(l)}$ (possibly after discarding a burn-in period), or to some Monte Carlo estimate of the posterior mode. Step 1 requires sampling from the posterior of the model $\bm{\gamma}$ given hyper-parameters $\bm{\omega}=\bm{\omega}^{(l-1)}$. This step has the same complexity as sampling $\bm{\gamma}$ in a fully Bayesian setting where no data integration is performed. Step 1 is easy to implement as long as one can compute or approximate $p(\bm{y}\mid\bm{\gamma})$, as is the case for generalized linear and generalized additive models, for example (see the opening of Section 5 for some discussion). Step 2 is typically also easy. First, note that $\bm{\omega}$ is often low-dimensional and that $\nabla_{\bm{\omega}}\log\pi(\bm{\gamma}\mid\bm{\omega})$ is usually easy to obtain (else, one can use automatic differentiation), enabling the use of gradient-based methods such as Hamiltonian Monte Carlo. Second, if $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j}\mbox{Bern}(\gamma_{j};m_{j}(\bm{\omega}))$ where $m_{j}$ is the inverse logit (or some other suitable) function, Step 2 requires sampling from the posterior of a logistic regression model where the observed outcomes are $\gamma_{1}^{(l)},\ldots,\gamma_{p}^{(l)}$, for which a pletora of algorithms are available.
@@
If marginal likelihoods $p(\bm{y}\mid\bm{\gamma})$ in Step 1 are hard to approximate, one could alternatively sample $\bm{\theta}^{(l)}\sim\pi(\bm{\theta}\mid\bm{y})$ using PDMPs such as the sticky Zig-Zag sampler (Bierkens et al., 2023), and subsequently obtain $\bm{\gamma}^{(l)}$ as a deterministic function of $\bm{\theta}^{(l)}$. Briefly, PDMPs only require being able to evaluate the likelihood $\pi(\bm{y}\mid\bm{\theta})$, and can handle situations where $\pi(\bm{\theta}\mid\bm{\omega})$ is such that some singletons have non-zero prior probability. For example, in variable selection we have non-zero $\pi(\theta_{j}=0\mid\bm{\omega})$ and, after sampling $\bm{\theta}^{(l)}\sim\pi(\bm{\theta}\mid\bm{y},\bm{\omega})$, one simply computes $\gamma_{j}^{(l)}=\mbox{I}(\theta_{j}^{(l)}\neq 0)$.
@@
If one only seeks a point estimate $\hat{\bm{\omega}}$, there’s yet another strategy. One can replace the model selection prior $\pi(\bm{\theta}\mid\bm{\omega})$, which places positive probability to zero parameter values, by a continuous prior that sets high probability close to zero (George and McCulloch, 1993). Typically this is achieved by replacing Dirac measures at 0 by Gaussian or Laplace distributions that have zero mean and small variance. In this setting it is possible to devise an EM algorithm to find a local mode of $\pi(\bm{\theta},\bm{\omega}\mid\bm{y})$ (Rockova and George, 2014; Avalos-Pacheco et al., 2022). Briefly, the algorithm views $\bm{\gamma}$ as a latent variable and defines a complete-data likelihood $p(\bm{y},\bm{\gamma}\mid\bm{\theta},\bm{\omega})$, similarly to the EM algorithm outlined in Section 5.2, except that the parameters are now $(\bm{\theta},\bm{\omega})$ rather than only $\bm{\omega}$. This strategy can be very fast computationally, because the EM updates do not involve integrals over $\bm{\gamma}$ and they may be even available in closed-form. Alas, the value of $\bm{\omega}$ featuring in the mode of $\pi(\bm{\theta},\bm{\omega}\mid\bm{y})$ is different from the mode of $\pi(\bm{\omega}\mid\bm{y})$. Whereas the latter has been well-studied theoretically (see Section 2), the former lacks theoretical guarantees (to our knowledge). Unless such guarantees are provided, we do not recommend this strategy.
@@ 6. Examples
To assess the improvements in inference brought by data integration, we compared the performance of our empirical Bayes framework relative to an analogous framework where $\bm{Z}$ only contains the intercept. We also considered a fully Bayesian counterpart that sets a Beta-Binomial model prior (Scott and Berger, 2010), which in our experience is an excellent default, and three penalized likelihood methods. The latter are the LASSO (Tibshirani, 1996), adaptive LASSO (ALASSO, Zou (2006)) and SCAD (Fan and Li, 2001), and we set their penalization parameters with 10-fold cross-validation using functions mylars and ncvreg in R packages parcor 0.2.6 and ncvreg 3.15.0 (respectively).
For our empirical Bayes methodology we used the EM algorithm from Section 5, which we implemented in function modelSelection_eBayes in R package mombf, using 1,000 full Gibbs iterations (i.e. $1000p$ model updates) to estimate log-gradients and hessians at each hyper-parameter value $\bm{\omega}$. For fully Bayesian inference under the Beta-Binomial prior we used function modelSelection, also at mombf. In both cases, we used 5,000 full Gibbs iterations to explore the model space. Regarding the prior on the coefficients, we set a product MOM prior with default prior dispersion ($g=1/3$, after the columns in $\bm{X}$ are standardized to unit sample variance) (Johnson and Rossell, 2012; Rossell and Telesca, 2017). As a sensitivity analysis, we also obtained results using Zellner’s unit information prior. For clarity of the figures and space constraints, Zellner’s prior results are in the supplement. The prior on the error variance was set to a (minimally informative) inverse gamma(0.01, 0.01). The R code to reproduce our analyses is at https://github.com/davidrusi/paper_examples/tree/main/2025_dataintegration_eBayes.
@@ 6.1. Simulation study
In each simulation scenario we considered three combinations of $n$ and $p$. The most challenging one features $(n=100,p=200)$, then we also considered $(n=200,p=200)$ and $(n=200,p=100)$. For each of these combinations, we generated 100 independent datasets and we report averaged results. For Bayesian methods, the point estimate was the Bayesian model averaging estimator. To obtain the frequentist power (proportion of truly active covariates that were selected) and false discovery rate (FDR, average proportion of false positives among the selected covariates), variable selection was based on their posterior inclusion probability being $\geq 0.95$.
Figure 3 displays the mean squared parameter estimation error (MSE) of the point estimates $\hat{\bm{\theta}}$, and the power. Figure 4 displays the FDR. To assess the potential benefits of data integration, the main comparisons of interest are our empirical Bayes framework vs. its counterpart where $\bm{Z}$ only features the intercept, and vs. the Beta-Binomial prior. To avoid cluttering the figures do not include the results for the intercept-only empirical Bayes. These are in the supplement, and show that the power and MSE are extremely similar to the Beta-Binomial and the FDR to empirical Bayes using all the meta-covariates. The main findings are that in Scenario 1 the empirical Bayes model prior attained significantly higher power and lower MSE than the Beta-Binomial (and the intercept-only empirical Bayes), and to a lesser extent also in Scenario 2. The improvements in power were marked in some cases, e.g. from 0.25 to 0.45 in Scenario 1 when $(n=100,p=200)$. The FDR was similar in most cases, except in Scenarios 2 and 3 where for the $(n=100,p=200)$ case the FDR was a higher for empirical Bayes. Overall, these results show that data integration can significantly improve inference when some of the meta-covariates are truly informative. In Scenario 3, where the meta-covariates were completely uninformative, the MSE and power of the empirical Bayes prior was similar to the Beta-Binomial, and the FDR was similar to that in Scenario 2. In Scenario 3 we do not expect empirical Bayes to help, but reassuringly it does not hurt much either. As mentioned, the results when using Zellner’s prior on the coefficients were qualitatively similar, see the supplementary results.
@@ 6.2. Colon cancer
In molecular biology it is common to use animal models (e.g. mice) to study genomic mechanisms such as gene expression. Findings from such studies often serve as a basis to study human diseases, for example to assess whether genes that played key roles in the animal model are also important in humans. Here we assess the performance of our empirical Bayes methodology in such a transfer learning task (from animal models to humans) in colon cancer.
@@
Specifically, Calon et al. (2012) compared mice where gene TGFB was silenced to mice where it was not, found a list of 172 genes that were associated to TGFB, and showed that TGFB plays a crucial role in colon cancer progression. We consider a colon cancer dataset analyzed in Rossell and Telesca (2017), which measures the expression of TGFB plus $p=1,000$ genes (including the 172 genes from the mouse list) for $n=262$ human patients. Both TGFB and the other genes were standardized to zero mean and unit variance. Given the importance of TGFB in colon cancer, it is of interest to understand how its expression is associated to other genes. To this end, we used our methodology to regress TGFB on the 1,000 genes, using as a meta-covariate the binary indicator of whether the gene was in the mouse shortlist or not. Hence, $\bm{Z}$ is a $1000\times 2$ matrix containing the intercept and said indicator (to be specific, coded with a sum-to-zero constraint, so that it is orthogonal to the intercept).
@@
As a purely exploratory analysis, TGFB showed stronger correlations with genes inside the mouse shortlist than with genes outside, see Figure 5 (left). While this suggests that the mouse findings may have value for humans, these correlations only measure marginal association, and are not concluding evidence that the mouse study helped identify genes that are predictive of TGFB (i.e., conditionally associated with TGFB, given all other genes). To assess this rigorously, we performed Bayesian variable selection with our empirical Bayes framework. For comparison, we repeated the exercise under a Beta-Binomial prior, which does not use the mouse data. We report results when using the pMOM prior on parameters, the results for Zellner’s prior are almost identical and are in the supplement. The estimated empirical Bayes prior inclusion probabilities were 0.032 for genes in the mouse list and 0.015 for those outside. Figure 5 (right) shows the posterior inclusion probabilities (PIPs) for all genes using empirical Bayes and using the Beta-Binomial. While the PIP for most genes was close to 0 under both, there are non-negligible differences for genes with large PIPs. Table 1 shows genes with PIP>>0.5 under either analysis. Empirical Bayes assigned higher PIPs to all these genes, e.g. 2 had PIP>>0.9 and 5 had PIP>>0.5, whereas for the Beta-Binomial there were 0 and 4 genes respectively.
@@
These differences in PIP are not huge, but they can have practical implications. For example, if one sets a stringent PIP threshold such as 0.95, only empirical Bayes would find a gene associated to TGFB. This gene is CILP, a cartilage intermediate layer protein that has been independently found to be associated to colon cancer progression (Wang et al., 2023). At a 0.9 threshold a second gene emerges, GAS1. This is a growth-arrest specific gene that is associated to colon cancer metastatis (Li et al., 2016). Note that both CILP and GAS1 were in the mouse shortlist, as well as HIC1 and ESM1, but not KCNJ5-AS1.

### v2

@@ 5.2. Expectation-Maximization algorithm
Consider an EM algorithm setting where the likelihood for the full data $(\bm{y},\bm{\gamma})$ is given by $p(\bm{y},\bm{\gamma}\mid\bm{\omega})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma}\mid\bm{\omega})$, and where $\bm{\gamma}$ are the latent variables in the usual definition of the EM algorithm. The algorithm requires an E-step where one obtains the expected full data log-likelihood with respect to $\bm{\gamma}$, given the data $\bm{y}$ and the current parameter estimate $\hat{\bm{\omega}}^{(k)}$. This requires obtaining $\pi(\bm{\gamma}\mid\bm{y},\hat{\bm{\omega}}^{(k)})$ and, if an exact computation is too costly (one cannot enumerate all models), one may use a stochastic EM algorithm where $\pi(\bm{\gamma}\mid\bm{y},\hat{\bm{\omega}}^{(k)})$ is replaced by the frequency of posterior samples visiting model $\bm{\gamma}$. Given such $\pi(\bm{\gamma}\mid\bm{y},\hat{\bm{\omega}}^{(k)})$, the M-step seeks to maximize
@@
$\nabla_{\bm{\omega}}f(\bm{\omega})=\sum_{j=1}^{p}\frac{\nabla_{\bm{\omega}}m_{j}(\bm{\omega})}{m_{j}(\bm{\omega})[1-m_{j}(\bm{\omega})]}\left[\pi(\gamma_{j}=1\mid\bm{y},\hat{\bm{\omega}}^{(k)})-\pi(\gamma_{j}=1\mid\bm{\omega})\right].$
In the particular case $m_{j}(\bm{\omega})=1/(1+e^{\bm{z}_{j}^{T}\bm{\omega}})$, then $\nabla_{\bm{\omega}}f(\bm{\omega})=\bm{Z}^{T}[\hat{\bm{\pi}}-{\bf m}(\bm{\omega})]$ and $\nabla_{\bm{\omega}}^{2}f(\bm{\omega})=-\bm{Z}^{T}\bm{D}(\bm{\omega})\bm{Z}$ where $\hat{\bm{\pi}}=(\pi(\gamma_{1}\mid\bm{y},\hat{\bm{\omega}}^{(k)}),\ldots,\pi(\gamma_{p}\mid\bm{y},\hat{\bm{\omega}}^{(k)}))^{T}$, ${\bf m}(\bm{\omega})=(m_{1}(\bm{\omega}),\ldots,m_{p}(\bm{\omega}))^{T}$ and $\bm{D}(\bm{\omega})=\mbox{diag}(m_{1}(\bm{\omega}),\ldots,m_{p}(\bm{\omega}))$.
Briefly, each M-step requires solving $\nabla_{\bm{\omega}}f(\bm{\omega})=\bm{A}[\hat{\bm{\pi}}-{\bf m}(\bm{\omega})]={\bf 0}$, where $\bm{A}$ is the $q\times p$ matrix with $j^{th}$ column equal to $\nabla_{\bm{\omega}}m_{j}(\bm{\omega})]/(m_{j}(\bm{\omega})[1-m_{j}(\bm{\omega})])$, ${\bf m}(\bm{\omega})\in\mathbb{R}^{p}$ the vector with $j^{th}$ entry equal to $m_{j}(\bm{\omega})$, and $\hat{\bm{\pi}}\in\mathbb{R}^{p}$ the vector with $j^{th}$ entry $\pi(\gamma_{j}=1\mid\bm{y},\hat{\bm{\omega}}^{(k)})$. One may use any standard optimization algorithm, and in our examples we used Newton-Raphson. When $m_{j}(\bm{\omega})$ is the inverse logit function, it is easy to see that the objective function of the M-step is strictly concave and therefore the maximum, if it exists, is unique.
@@
Let $\bm{U}$ be the sub-matrix containing the unique rows of the $p\times q$ matrix $\bm{Z}$. Suppose that $\bm{U}$ has $q$ rows and rank $q$, $m_{j}(\bm{\omega})=1/(1+e^{-\bm{z}_{j}^{T}\bm{\omega}})$, and $\bm{\omega}\sim N({\bf 0},g(\bm{Z}^{T}\bm{Z}/p)^{-1})$. Then the posterior mode $\hat{\bm{\omega}}$ satisfies $\nabla_{\bm{\omega}}f(\hat{\bm{\omega}})+\nabla_{\bm{\omega}}\log\pi(\hat{\bm{\omega}})={\bf 0}$, where $\hat{\bm{\omega}}=\bm{U}^{-1}\tilde{\bm{\omega}}$ and $\tilde{\omega}_{j}=h(a_{j},g_{\omega}p)$, where $h(a,c)$ is the solution to $1/(1+e^{-w})+w/c=a$, $a_{j}=\sum_{i=1}^{p}\tilde{z}_{ij}^{T}\hat{\pi}_{i}/\sum_{i=1}^{p}z_{ij}$, $\tilde{\bm{Z}}=\bm{Z}\bm{U}^{-1}$ and $\hat{\bm{\pi}}$ is as in Corollary 1.
@@ 5.3. Other computational methods
Upon convergence, $\bm{\omega}^{(l)}$ is by definition a sample from $\pi(\bm{\omega}\mid\bm{y})$. One may set $\hat{\bm{\omega}}$ to the posterior mean $L^{-1}\sum_{l=1}^{L}\bm{\omega}^{(l)}$ (possibly after discarding a burn-in period), or to some Monte Carlo estimate of the posterior mode. Step 1 requires sampling from the posterior of the model $\bm{\gamma}$ given hyper-parameters $\bm{\omega}=\bm{\omega}^{(l-1)}$. This step has the same complexity as sampling $\bm{\gamma}$ in a fully Bayesian setting where no data integration is performed. Step 1 is easy to implement as long as one can compute or approximate $p(\bm{y}\mid\bm{\gamma})$, as is the case for generalized linear and generalized additive models, for example (see the opening of Section 5 for some discussion). Step 2 is typically also easy. First, note that $\bm{\omega}$ is often low-dimensional and that $\nabla_{\bm{\omega}}\log\pi(\bm{\gamma}\mid\bm{\omega})$ is usually easy to obtain (else, one can use automatic differentiation), enabling the use of gradient-based methods such as Hamiltonian Monte Carlo. Second, if $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j}\mbox{Bern}(\gamma_{j};m_{j}(\bm{\omega}))$ where $m_{j}$ is the inverse logit (or some other suitable) function, Step 2 requires sampling from the posterior of a logistic regression model where the observed outcomes are $\gamma_{1}^{(l)},\ldots,\gamma_{p}^{(l)}$, for which a plethora of algorithms are available.
@@
If marginal likelihoods $p(\bm{y}\mid\bm{\gamma})$ in Step 1 are hard to approximate, one could alternatively sample $\bm{\theta}^{(l)}\sim\pi(\bm{\theta}\mid\bm{y})$ using PDMPs such as the sticky Zig-Zag sampler (Bierkens et al., 2023), and subsequently obtain $\bm{\gamma}^{(l)}$ as a deterministic function of $\bm{\theta}^{(l)}$. Briefly, PDMPs only require being able to evaluate the likelihood $p(\bm{y}\mid\bm{\theta})$, and can handle situations where $\pi(\bm{\theta}\mid\bm{\omega})$ is such that some singletons have non-zero prior probability. For example, in variable selection we have non-zero $\pi(\theta_{j}=0\mid\bm{\omega})$ and, after sampling $\bm{\theta}^{(l)}\sim\pi(\bm{\theta}\mid\bm{y},\bm{\omega})$, one simply computes $\gamma_{j}^{(l)}=\mbox{I}(\theta_{j}^{(l)}\neq 0)$.
@@
If one only seeks a point estimate $\hat{\bm{\omega}}$, there’s yet another strategy. One can replace the model selection prior $\pi(\bm{\theta}\mid\bm{\omega})$, which places positive probability to zero parameter values, by a continuous prior that sets high probability close to zero (George and McCulloch, 1993). Typically, this is achieved by replacing Dirac measures at 0 by Gaussian or Laplace distributions that have zero mean and small variance. In this setting it is possible to devise an EM algorithm to find a local mode of $\pi(\bm{\theta},\bm{\omega}\mid\bm{y})$ (Rockova and George, 2014; Avalos-Pacheco et al., 2022). Briefly, the algorithm views $\bm{\gamma}$ as a latent variable and defines a complete-data likelihood $p(\bm{y},\bm{\gamma}\mid\bm{\theta},\bm{\omega})$, similarly to the EM algorithm outlined in Section 5.2, except that the parameters are now $(\bm{\theta},\bm{\omega})$ rather than only $\bm{\omega}$. This strategy can be very fast computationally, because the EM updates do not involve integrals over $\bm{\gamma}$ and they may be even available in closed-form. However, the value of $\bm{\omega}$ featuring in the mode of $\pi(\bm{\theta},\bm{\omega}\mid\bm{y})$ is different from the mode of $\pi(\bm{\omega}\mid\bm{y})$. Whereas the latter has been well-studied theoretically (see Section 2), the former lacks theoretical guarantees (to our knowledge). Unless such guarantees are provided, we do not recommend this strategy.
@@ 6. Examples
To assess the improvements in inference brought by data integration, we compared the performance of our empirical Bayes framework relative to an analogous framework where $\bm{Z}$ only contains the intercept. We also considered the Beta-Binomial prior (Scott and Berger, 2010) as a fully Bayesian counterpart to our framework where $\bm{Z}$ only contains the intercept, as in our experience this method is an excellent default, and three penalized likelihood methods. The latter are the LASSO (Tibshirani, 1996), adaptive LASSO (ALASSO, Zou (2006)) and SCAD (Fan and Li, 2001), and we set their penalization parameters with 10-fold cross-validation using functions mylars and ncvreg in R packages parcor 0.2.6 and ncvreg 3.15.0 (respectively).
For our empirical Bayes methodology we used the EM algorithm from Section 5, which we implemented in function modelSelection_eBayes in R package modelSelection, using $M=1,000$ full Gibbs iterations to estimate log-gradients and hessians at each hyper-parameter value $\bm{\omega}$. For fully Bayesian inference under the Beta-Binomial prior we used function modelSelection, also at modelSelection. In both cases, we used $L=5,000$ full Gibbs iterations to explore the model space. Appendix 12 shows that very similar results are obtained with $M=100$ and $L=1,000$. It also shows that run times are comparable to those under the Beta-Binomial prior. Regarding the prior on the coefficients, we set a product MOM prior with default prior dispersion ($g=1/3$, after the columns in $\bm{X}$ are standardized to unit sample variance) (Johnson and Rossell, 2012; Rossell and Telesca, 2017). As a sensitivity analysis, we also obtained results using Zellner’s prior with the default $g=1$, and also for $g\in\{0.01,10\}$. For clarity of the figures and space constraints, Zellner’s prior results are in the supplement. The prior on the error variance $\phi$ was set to a (minimally informative) inverse gamma(0.01, 0.01). The R code to reproduce our analyses is at https://github.com/davidrusi/paper_examples/tree/main/2025_dataintegration_eBayes.
@@ 6.1. Simulation study
In each simulation scenario we considered three combinations of $n$ and $p$. The most challenging one features $(n=100,p=200)$, then we also considered $(n=200,p=200)$ and $(n=200,p=100)$. For each of these combinations, we generated 100 independent datasets, and we report averaged results. For Bayesian methods, the point estimate was the Bayesian model averaging estimator. To obtain the frequentist power (proportion of truly active covariates that were selected) and false discovery rate (FDR, average proportion of false positives among the selected covariates), variable selection was based on their posterior inclusion probability being $\geq 0.95$.
Figure 3 displays the mean squared parameter estimation error (MSE) of the point estimates $\hat{\bm{\theta}}$, and the power. Figure 4 displays the FDR. To assess the potential benefits of data integration, the main comparisons of interest are our empirical Bayes framework vs. its counterpart where $\bm{Z}$ only features the intercept, and vs. the Beta-Binomial prior. To avoid cluttering the figures, we do not include the results for the intercept-only empirical Bayes. These are in the supplement, and show that the power and MSE are extremely similar to the Beta-Binomial and the FDR to empirical Bayes using all the meta-covariates. The main findings are that in Scenario 1 the empirical Bayes model prior attained significantly higher power and lower MSE than the Beta-Binomial (and the intercept-only empirical Bayes), and to a lesser extent also in Scenario 2. The improvements in power were marked in some cases, e.g. from 0.25 to 0.45 in Scenario 1 when $(n=100,p=200)$. The FDR was similar in most cases, except in Scenarios 2 and 3 where for the $(n=100,p=200)$ case the FDR was a higher for empirical Bayes. Overall, these results show that data integration can significantly improve inference when some of the meta-covariates are truly informative. In Scenario 3, where the meta-covariates were completely uninformative, the MSE and power of the empirical Bayes prior was similar to the Beta-Binomial, and the FDR was similar to that in Scenario 2. In Scenario 3 we do not expect empirical Bayes to help, but reassuringly it does not hurt much either. As mentioned, the results when using Zellner’s prior on the coefficients were qualitatively similar, see the supplementary results.
@@ 6.2. Colon cancer
In molecular biology it is common to use animal models (e.g. mice) to study genomic mechanisms such as gene expression. Findings from such studies often serve as a basis to study human diseases, for example to assess whether genes that played key roles in the animal model are also important in humans. We assess our empirical Bayes methodology in such a transfer learning task (from mice to humans) in colon cancer.
@@
Specifically, Calon et al. (2012) studied a list of 172 genes that were associated to TGFB in mice, and showed that TGFB plays a crucial role in colon cancer progression. We consider a colon cancer dataset analysed in Rossell and Telesca (2017), which measures the expression of TGFB plus $p=1,000$ genes (including the 172 genes from the mouse list) for $n=262$ human patients. Both TGFB and the other genes were standardized to zero mean and unit variance. Given the importance of TGFB in colon cancer, it is of interest to understand how its expression is associated with other genes. To this end, we used our methodology to regress TGFB on the 1,000 genes, using as a meta-covariate the binary indicator of whether the gene was in the mouse shortlist or not. Hence, $\bm{Z}$ is a $1000\times 2$ matrix containing the intercept and said indicator (to be specific, coded with a sum-to-zero constraint, so that it is orthogonal to the intercept).
@@
As a purely exploratory analysis, TGFB showed stronger correlations with the 172 genes inside the mouse shortlist than with other genes, see Figure 5 (left). While this suggests that the mouse findings may have value for humans, these correlations only measure marginal association, and are not concluding evidence that the mouse study helped identify genes that are predictive of TGFB (i.e., conditionally associated with TGFB, given all other genes). To assess this rigorously, we performed Bayesian variable selection with our empirical Bayes framework. For comparison, we repeated the exercise under a Beta-Binomial prior, which does not use the mouse data. The run time under the Beta-Binomial prior was 14 seconds and for our empirical Bayes framework it was 49 seconds (Ubuntu laptop, 13th gen Intel core i7, 32Gb RAM). We report results when using the pMOM prior on parameters, the results for Zellner’s prior are almost identical and are in the supplement. The estimated empirical Bayes prior inclusion probabilities were 0.032 for genes in the mouse list and 0.015 for those outside, supporting that genes identified in mice are more likely to be relevant in humans. Figure 5 (right) shows the posterior inclusion probabilities (PIPs) for all genes using empirical Bayes and using the Beta-Binomial. While the PIP for most genes was close to 0 under both, there are non-negligible differences for genes with large PIPs. Table 1 shows genes with PIP>>0.5 under either analysis. Empirical Bayes assigned higher PIPs to all these genes, e.g. 2 had PIP>>0.9 and 5 had PIP>>0.5, whereas for the Beta-Binomial there were 0 and 4 genes respectively.
@@
These differences in PIP are not huge, but they can have practical implications. For example, if one sets a stringent PIP threshold such as 0.95, only empirical Bayes would find a gene associated with TGFB. This gene is CILP, a cartilage intermediate layer protein that has been independently found to be associated with colon cancer progression (Wang et al., 2023). At a 0.9 threshold a second gene emerges, GAS1. This is a growth-arrest specific gene that is associated with colon cancer metastasis (Li et al., 2016). Note that both CILP and GAS1 were in the mouse shortlist, as well as HIC1 and ESM1, but not KCNJ5-AS1.